OpenStudy (anonymous):

Hey guys, can you help me out with this? Without graphing, is the system independent, dependent, or inconsistent? y=2x-8, 2x-y=8 .

OpenStudy (anonymous):

Oh hello again Viktor.ray lol..

OpenStudy (anonymous):

:o Oh my goodness so many people viewing my question yey (:

OpenStudy (anonymous):

Yup Hello Megan first eqtn: y=2x-8 secnd eqtn on rearranging: y=2x-8 Hence they are dependent (same equations lol)

OpenStudy (anonymous):

That is what I as thinking just wanted to make sure x.x Do you have any other examples of systems that would be independant or inconsistant?

OpenStudy (anonymous):

Well if the second equation had been y-2x=8 then it would have been an inconsistent system as then the 1st eq would give y=2x-8 and 2nd eq would give y=2x+8 so basically for the two to be equal, -8=+8 which is impossible. Hence that would be an inconsistent system.. Independent not recollecting anything right now off the top of my mind will send you some links in some time if you want

OpenStudy (anonymous):

That would be nice Since it seems most of the problems they give us end up being inconsistant :P

OpenStudy (anonymous):

lol if ur teachers/profs are consistent about giving inconsistent stuff, that just makes ur work easier now, dsnt it? :P ;)

OpenStudy (anonymous):

Bahaha I guess it kinda does x)